factorize 3p^2-2p-5
1
Expert's answer
2012-10-16T09:49:21-0400
we find roots of the equation

3p^2-2p-5=0
D=4+4*3*5=4*16=64
p1,2=(2+-8)/6 so p1=-1, p2=5/3
so
3p^2-2p-5=3(p+1)(p-5/3)